# PDRL: Post-hoc Descriptor-based Residual Learning for Uncertainty-Aware Machine Learning Potentials

### v1

@@ Abstract
Ensemble method is considered the gold standard for uncertainty quantification (UQ) for machine learning interatomic potentials (MLIPs). However, their high computational cost can limit its practicality. Alternative techniques, such as Monte Carlo dropout and deep kernel learning, have been proposed to improve computational efficiency; however, some of these methods cannot be applied to already trained models and may affect the prediction accuracy. In this paper, we propose a simple and efficient post-hoc framework for UQ that leverages the descriptor of a trained graph neural network potential to estimate residual errors. We refer to this method as post-hoc descriptor-based residual-based learning (PDRL). PDRL models the discrepancy between MLIP predictions and ground truth values, allowing these residuals to act as proxies for prediction uncertainty. We explore multiple variants of PDRL and benchmark them against established UQ methods, evaluating both their effectiveness and limitations.
@@ 1 Introduction
Machine learning interatomic potentials (MLIPs) are transforming materials science by enabling atomic-scale simulations with the accuracy of quantum mechanical methods but at orders-of-magnitude higher computational efficiency [Behler and Parrinello, 2007, Bartók et al., 2010, Shapeev, 2016, Unke et al., 2021, Behler, 2016, Deringer et al., 2019, Chanussot et al., 2021, Batatia et al., 2022, Takamoto et al., 2022, Batatia et al., 2023]. Despite their promise, the predictive reliability of MLIPs remains a critical concern, especially for atomic configurations outside the training data distribution [Wollschläger et al., 2023, Tan et al., 2023, Kellner and Ceriotti, 2024, Dai et al., 2025]. Robust uncertainty quantification (UQ) is essential for assessing model reliability, guiding decision-making, and ensuring trustworthy simulation outcomes.
@@
Several methods for UQ have been explored for MLIPs [Janet et al., 2019, Zhu et al., 2023, Vita et al., 2025]. Ensemble approaches, which aggregate predictions from multiple independently trained models, are widely regarded as the gold standard due to their effectiveness and their simplicity to implement [Lakshminarayanan et al., 2017, Tan et al., 2023]. However, ensembles are computationally expensive, particularly in settings with large-scale data or complex models like graph neural networks (GNN) [Batatia et al., 2022, Schütt et al., 2017, Gasteiger et al., 2020, Schütt et al., 2021]. Several alternative techniques have been proposed to address these limitations. Monte Carlo (MC) dropout [Gal and Ghahramani, 2016] suggests to enable dropout during inference to introduce stochasticity into predictions at test time, and deep kernel learning combines neural networks with Gaussian processes [Wollschläger et al., 2023]. While these approaches improve efficiency, many of them require modification of the training pipeline or are incompatible with post-hoc applications, limiting their flexibility in scenarios where models are pre-trained and fixed.
@@
There also exists another line of research that has explored UQ in MLIPs through the usage of model descriptors and prediction errors, based on the assumption that high prediction error structures should have higher uncertainty than low prediction error on average. Vita et al. [2025] proposed loss trajectory analysis for uncertainty (LTAU), which leverages per-atom force error predictions to train an uncertainty model. While effective, this method requires logging the loss trajectory for every atom and is limited to capturing uncertainty in per-atom force predictions. In Orb-v3 [Rhodes et al., 2025], prediction errors are discretized in a manner inspired by pLDDT in Alphafold [Jumper et al., 2021], enabling joint optimization of the UQ objective alongside the prediction objective during model training. In contrast to post-hoc methods, UQ objective of pLDDT is optimized jointly with the model training process. For methods that are post-hoc that only utilizes model descriptor features, Janet et al. [2019] proposed to calculate the average feature distance between a test point and its k-nearest neighbors in the training data using revised autocorrelation descriptors. Zhu et al. [2023] demonstrated the utility of fitting a Gaussian mixture model (GMM) for UQ in NequIP [Batzner et al., 2022]. To the best of our knowledge, there has been limited study of post-hoc methods that explicitly estimate prediction errors using only the trained model, without requiring detailed training logs or specialized optimization procedures during training.
@@
In this paper, we explore post-hoc descriptor-based residual learning (PDRL) that utilize model descriptors of a trained graph neural network potential to estimate prediction error. These residuals act as proxies for prediction error, enabling efficient and scalable UQ without modifying the original model architecture or training process. We investigate multiple variants of PDRL for energy and force errors: error-norm learning and deviation learning. Error-norm learning predicts a scalar representing the norm of the error, while deviation learning models the intrinsic error directly, maintaining the same format as the original prediction.
@@ 2 Notation
Let be a molecular or material structure in the input space . To learn an MLIP, it is common that we use a training dataset with molecular or meterial structures: . Each structure consists of atoms and is represented as . Here, represents the atomic position information in three-dimensional space, and encodes the atomic numbers corresponding to each atom within the structure. For example, indicates that the first atom in structure is hydrogen. ’s energy label and force label are also provided for each structure. With this training data, the goal of MLIP training is to accurately learn a real-valued function , which predicts energy given a structure of interest. Not only energy, we also expect that the force prediction is correctly predicted, where one can obtain the force information of atom given a predicted energy by calculating a negative gradient of (obtained via ) with respect to atomic position, i.e., . We note that some MLIPs directly predict forces (e.g., ForceNet [Hu et al., 2021]), but these are beyond the scope of this paper.
3 Post-hoc descriptor-based residual learning (PDRL)
Given structure , in general, we can extract its descriptor in graph neural networks. In this paper, we focus on a message-passing atomic cluster expansion (MACE) [Batatia et al., 2022] model, where its descriptor (aka. features) can be extracted. For , we denote a function , which maps a structure to a descriptor for each atom, where indicates the dimension of the descriptor. With slight abuse of notation, we denote a descriptor of atom . Given a structure along with its energy and force , , , we can calculate energy residual and force residual , where and indicate energy and force predictions of an MLIP. Given structures and a trained model, we can prepare the training data for PDRL: .
@@ 3.1 Energy residual learning
A naive approach for residual energy learning might involve designing a single model that processes an entire molecular structure with atoms and outputs the residual. However, this straightforward approach fails to naturally preserve permutational invariance and lacks the flexibility to handle systems of varying atomic sizes. To address these challenges, we train a multilayered perceptron designed to take as input a descriptor for a single atom and output its corresponding scalar value . Since the energy depends on the amount of substance, we model the energy residual as the sum of the atom-wise score function. This approach allows the model to operate atom-wise, ensuring invariance properties and scalability to arbitrarily-sized structures. In error norm learning, we can calculate a structure-wise squared loss by . For the energy deviation learning, we learn directly, i.e., .
@@ 3.2 Force residual learning
Unlike energy, we designed a model that directly predicts force residuals using descriptors of individual atoms. In error norm learning, we learn a function to estimate the Euclidean norm of the force error by minimizing the atomwise loss , where is a real-valued function similarly to energy residual learning and denotes the euclidean norm. In deviation learning, we minimize the atomwise loss , where is a vector-valued function. Similarly to energy residual learning, we use a multilayered perceptron to model the force residual function and .
@@ 4 Experimental results
[…]
| MC-dropout [Gal and Ghahramani, 2016] | -0.02 (0.02) | -0.02 (0.02) | -0.03 (0.02) | -0.05 (0.05) | 0.20 (0.03) | |
| GMM [Zhu et al., 2023] | 0.07 (0.05) | 0.07 (0.09) | 0.13 (0.08) | 0.64 (0.05) | 0.06 (0.03) | |
| kNN [Janet et al., 2019] | 0.06 (0.04) | 0.06 (0.07) | 0.09 (0.06) | 0.64 (0.05) | -0.05 (0.03) | |
| PDRL-norm | 0.12 (0.13) | -0.01 (0.04) | -0.09 (0.03) | 0.62 (0.07) | 0.30 (0.03) | |
| PDRL-diff | 0.18 (0.14) | 0.16 (0.16) | 0.21 (0.14) | 0.87 (0.03) | 0.26 (0.02) | |
| MC-dropout [Gal and Ghahramani, 2016] | 0.24 (0.03) | 0.27 (0.02) | 0.27 (0.06) | 0.87 (0.01) | 0.68 (0.01) | |
| GMM [Zhu et al., 2023] | 0.58 (0.01) | 0.67 (0.03) | 0.68 (0.02) | 0.96 (0.01) | 0.64 (0.04) | |
| kNN [Janet et al., 2019] | 0.52 (0.02) | 0.61 (0.04) | 0.65 (0.02) | 0.96 (0.01) | 0.54 (0.01) | |
| PDRL-norm | 0.67 (0.02) | 0.71 (0.04) | 0.69 (0.02) | 0.98 (0.00) | 0.92 (0.00) | |
| PDRL-diff | 0.53 (0.02) | 0.58 (0.04) | 0.57 (0.01) | 0.96 (0.01) | 0.85 (0.01) |
[…]
4.1 Uncertainty-error correlation evaluation
Table 1 reports the 5-trial average Spearman correlation for each method across datasets. The results indicate that PDRL-diff performs relatively well in predicting energy uncertainties, whereas PDRL-norm is more effective for force uncertainties. Notably, the performance gap between GMM and kNN is larger in HME21. For readers who are interested, we provide an additional analysis using principal component analysis of HME21 in Appendix E.
[…]
Table 2 summarizes the performance of various OOD detection methods. Ensemble, kNN, GMM, and PDRL-diff consistently achieve strong OOD detection performance across the benchmark. In comparison, PDRL-norm performs less effectively, while MC-dropout yields the lowest performance in terms of both Spearman correlation and AUC. Although PDRL-norm achieves the best performance in Table 1, it proves less effective in the OOD setting in our experiments.
| Method | High temp. | Hexagonal | Cubic | Swap | All | |||||
| Corr. | AUC | Corr. | AUC | Corr. | AUC | Corr. | AUC | Corr. | AUC | |
| Ensemble | 0.98 | 1.00 | 0.88 | 0.94 | 0.95 | 1.00 | 0.76 | 1.00 | 0.90 | 0.99 |
| MC-dropout [Gal and Ghahramani, 2016] | 0.92 | 1.00 | 0.54 | 0.63 | 0.81 | 0.84 | 0.62 | 0.82 | 0.72 | 0.82 |
| GMM [Zhu et al., 2023] | 0.98 | 1.00 | 0.74 | 1.00 | 0.73 | 1.00 | 0.77 | 1.00 | 0.81 | 1.00 |
| kNN [Janet et al., 2019] | 0.98 | 1.00 | 0.80 | 0.99 | 0.75 | 1.00 | 0.76 | 1.00 | 0.82 | 1.00 |
| PDRL-norm | 0.99 | 1.00 | 0.82 | 0.82 | 0.78 | 0.82 | 0.70 | 0.99 | 0.82 | 0.91 |
| PDRL-diff | 0.97 | 1.00 | 0.89 | 0.97 | 0.80 | 1.00 | 0.81 | 1.00 | 0.87 | 0.99 |
5 Discussions
When PDRL outperform KNN, GMM? While descriptor-based methods perform competitively in our experiments, PDRL clearly outperforms KNN and GMM in error–uncertainty correlation on the HME21 dataset. We hypothesize that when a dataset contains many elements, descriptor information alone may be difficult to capture the error correlation, and incorporating the prediction error signal can enhance the uncertainty estimation by aligning it more closely with the true prediction error.
Which approach is better: error norm learning or deviation learning? In Table 1, deviation learning outperforms error-norm learning for energy Spearman correlation, whereas the opposite trend is observed for force correlation. This difference likely arises from the nature of the targets: energy deviation is a scalar, where retaining the error sign aids learning, while force deviation is a three-dimensional vector, making direct estimation more challenging. Using the force error norm reduces this complexity, improving Spearman correlation. However, as shown in Table 2, using force uncertainty, error-norm learning underperforms deviation learning in OOD detection, indicating the need for further investigation into the advantages and limitations of these approaches. We hypothesize that compressing errors into a norm may be detrimental for OOD detection, while remaining sufficiently informative for in-distribution predictions.
Conclusion and future work We studied the effectiveness of post-hoc descriptor-based residual learning (PDRL) for uncertainty estimation in machine learning interatomic potentials (MLIPs). PDRL achieves superior Spearman correlation with prediction errors compared to other methods. However, PDRL-norm underperforms in OOD detection relative to GMM and KNN descriptor-based approaches, highlighting its limitations. Future work will explore extending PDRL to active learning pipelines and other MLIP applications, further assessing its versatility, as well as investigating improved training methods to enhance PDRL’s performance.
Appendix A Broader Impact
This work focuses on developing uncertainty estimation techniques called post-hoc descriptor-based residual learning (PDRL) for machine learning interatomic potentials. PDRL is designed to improve the reliability and robustness of simulations in chemistry, materials science, and related fields. Importantly, our research does not involve human subjects, personal data, or any form of unethical experimentation. The techniques we propose are purely computational and are evaluated on standard benchmark datasets and simulated molecular systems.
Appendix B Dataset statistics
Table 3 shows statistics of all datasets used in this paper.
[…]
@@ Appendix B Dataset statistics
Appendix C Training details and hyperparameters

### v2

# Machine Learning and the Physical Sciences (ML4PS) P-DRUM: Post-hoc Descriptor-based Residual Uncertainty Modeling for Machine Learning Potentials

@@ Abstract
Ensemble method is considered the gold standard for uncertainty quantification (UQ) in machine learning interatomic potentials (MLIPs). However, their high computational cost can limit its practicality. Alternative techniques, such as Monte Carlo dropout and deep kernel learning, have been proposed to improve computational efficiency; however, some of these methods cannot be applied to already trained models and may affect the prediction accuracy. In this paper, we propose a simple and efficient post-hoc framework for UQ that leverages the descriptor of a trained graph neural network potential to estimate residual errors. We refer to this method as post-hoc descriptor-based residual uncertainty modeling (P-DRUM). P-DRUM models the discrepancy between MLIP predictions and ground truth values, allowing these residuals to act as proxies for prediction uncertainty. We explore multiple variants of P-DRUM and benchmark them against established UQ methods, evaluating both their effectiveness and limitations.
@@ 1 Introduction
Machine learning interatomic potentials (MLIPs) are transforming materials science by enabling atomic-scale simulations with the accuracy of quantum mechanical methods but at orders-of-magnitude higher computational efficiency [bpnet, gap, mtp, unke2021machine, behler2016perspective, deringer2019machine, oc20, batatia2022mace, takamoto2022towards, batatia2023foundation]. Despite their promise, the predictive reliability of MLIPs remains a critical concern, especially for atomic configurations outside the training data distribution [wollschlager2023uncertainty, tan2023single, kellner2024uncertainty, dai2025uncertainty]. Robust uncertainty quantification (UQ) is essential for assessing model reliability, guiding decision-making, and ensuring trustworthy simulation outcomes.
@@
Several methods for UQ have been explored for MLIPs [janet2019quantitative, zhu2023fast, vita2025ltau]. Ensemble approaches, which aggregate predictions from multiple independently trained models, are widely regarded as the gold standard due to their effectiveness and their simplicity to implement [lakshminarayanan2017simple, tan2023single]. However, ensembles are computationally expensive, particularly in settings with large-scale data or complex models like graph neural networks (GNN) [batatia2022mace, schutt2017schnet, dimenet, schutt2021equivariant]. Several alternative techniques have been proposed to address these limitations. Monte Carlo (MC) dropout [gal2016dropout] suggests to enable dropout during inference to introduce stochasticity into predictions at test time, and deep kernel learning combines neural networks with Gaussian processes [wollschlager2023uncertainty, rajput2023uncertainty, liu2020simple]. While these approaches improve efficiency, many of them require modification of the training pipeline or are incompatible with post-hoc applications, limiting their flexibility in scenarios where models are pre-trained and fixed.
@@
Model descriptors have been shown to be effective in various downstream applications such as chemical property prediction [descriptor1, descriptor2]. Also, they have been combined with prediction errors for UQ, under the assumption that structures with higher prediction errors tend to exhibit greater uncertainty on average. vita2025ltau proposed loss trajectory analysis for uncertainty (LTAU), which leverages per-atom force error predictions to train an uncertainty model. While effective, this method requires logging the loss trajectory for every atom and is limited to capturing uncertainty in per-atom force predictions. In Orb-v3 [rhodes2025orb], prediction errors are discretized in a manner inspired by pLDDT in Alphafold [alphafold], enabling joint optimization of the UQ objective alongside the prediction objective during model training. In contrast to post-hoc methods, UQ objective of pLDDT is optimized jointly with the model training process. For methods that are post-hoc that only utilizes model descriptor features, janet2019quantitative proposed to calculate the average feature distance between a test point and its k-nearest neighbors in the training data using revised autocorrelation descriptors. zhu2023fast demonstrated the utility of fitting a Gaussian mixture model (GMM) for UQ in NequIP [nequip]. To the best of our knowledge, there has been limited study of post-hoc methods that explicitly estimate prediction errors using only the trained model, without requiring detailed training logs or specialized optimization procedures during training.
@@
In this paper, we explore post-hoc descriptor-based residual uncertainty modeling (P-DRUM) that utilize model descriptors of a trained graph neural network potential to estimate prediction error. These residuals act as proxies for prediction error, enabling efficient and scalable UQ without modifying the original model architecture or training process. We investigate multiple variants of P-DRUM for energy and force errors: error-norm learning and deviation learning. Error-norm learning predicts a scalar representing the norm of the error, while deviation learning models the intrinsic error directly, maintaining the same format as the original prediction.
@@ 2 Notation
Let be a molecular or material structure in the input space . To learn an MLIP, it is common that we use a training dataset with molecular or meterial structures: . Each structure consists of atoms and is represented as . Here, represents the atomic position information in three-dimensional space, and encodes the atomic numbers corresponding to each atom within the structure. For example, indicates that the first atom in structure is hydrogen. ’s energy label and force label are also provided for each structure. With this training data, the goal of MLIP training is to accurately learn a real-valued function , which predicts energy given a structure of interest. Not only energy, we also expect that the force prediction is correctly predicted, where one can obtain the force information of atom given a predicted energy by calculating a negative gradient of (obtained via ) with respect to atomic position. We note that some MLIPs directly predict forces (e.g., ForceNet [hu2021forcenet]), but these are beyond the scope of this paper.
3 Post-hoc descriptor-based residual uncertainty modeling (P-DRUM)
Given structure , in general, we can extract its descriptor in graph neural networks. In this paper, we focus on a message-passing atomic cluster expansion (MACE) [batatia2022mace] model, where its descriptor (aka. features) can be extracted. For , we denote a function , which maps a structure to a descriptor for each atom, where indicates the dimension of the descriptor. we denote a descriptor of atom in structure . Given a structure along with its energy and force , , , we can calculate energy residual and force residual , where and indicate energy and force predictions of an MLIP. Given structures and a trained model, we can prepare the training data for P-DRUM: . Using the residual training data , we train a supervised model based on a multilayer perceptron (MLP) architecture, which offers significantly faster inference than MLIP models such as MACE. Figure 1 shows the overview of P-DRUM.
@@ 3.1 Energy residual learning
A naive approach for residual energy learning might involve designing a single model that processes an entire molecular structure with atoms and outputs the residual. However, this straightforward approach fails to naturally preserve permutational invariance and lacks the flexibility to handle systems of varying atomic sizes. To address these challenges, we train a multilayered perceptron designed to take as input a descriptor for a single atom and output its corresponding scalar value . Since the energy depends on the whole structure, but each structure can have different number of atoms, we model the energy residual as the sum of the atom-wise score function. This approach allows the model to operate atom-wise, ensuring invariance properties and scalability to arbitrarily-sized structures. In error norm learning, we can calculate a structure-wise squared loss by . For the energy deviation learning, we learn directly, i.e., .
@@ 3.2 Force residual learning
Unlike energy, we designed a model that directly predicts force residuals using descriptors of individual atoms. In error norm learning, we learn a function to estimate the Euclidean norm of the force error by minimizing the atomwise loss , where is a real-valued function similarly to energy residual learning and denotes the euclidean norm. In deviation learning, for each atom in a structure, we minimize the average of the coordinate-wise force-error loss: , where is a vector-valued function. Uncertainty can be calculated by the Euclidean norm of the -dimensional output of . Similarly to energy residual learning, we use a multilayered perceptron to model the force residual function.
@@ 4 Experimental results
| rMD17 [rmd17] | HME21 [takamoto2022towards] | ||||
@@
| Error type | Method | Uracil | Salicylic | Malondi- | HME21 | |
| aldehyde | ||||||
| MC-dropout [gal2016dropout] | -0.02 (0.02) | -0.02 (0.02) | -0.03 (0.02) | -0.05 (0.05) | 0.20 (0.03) | |
| GMM [zhu2023fast] | 0.07 (0.05) | 0.07 (0.09) | 0.13 (0.08) | 0.64 (0.05) | 0.06 (0.03) | |
| kNN [janet2019quantitative] | 0.06 (0.04) | 0.06 (0.07) | 0.09 (0.06) | 0.64 (0.05) | -0.05 (0.03) | |
| P-DRUM-norm | 0.12 (0.13) | -0.01 (0.04) | -0.09 (0.03) | 0.62 (0.07) | 0.30 (0.03) | |
| P-DRUM-diff | 0.18 (0.14) | 0.16 (0.16) | 0.21 (0.14) | 0.87 (0.03) | 0.26 (0.02) | |
| MC-dropout [gal2016dropout] | 0.24 (0.03) | 0.27 (0.02) | 0.27 (0.06) | 0.87 (0.01) | 0.68 (0.01) | |
| GMM [zhu2023fast] | 0.58 (0.01) | 0.67 (0.03) | 0.68 (0.02) | 0.96 (0.01) | 0.64 (0.04) | |
| kNN [janet2019quantitative] | 0.52 (0.02) | 0.61 (0.04) | 0.65 (0.02) | 0.96 (0.01) | 0.54 (0.01) | |
| P-DRUM-norm | 0.67 (0.02) | 0.71 (0.04) | 0.69 (0.02) | 0.98 (0.00) | 0.92 (0.00) | |
| P-DRUM-diff | 0.53 (0.02) | 0.58 (0.04) | 0.57 (0.01) | 0.96 (0.01) | 0.85 (0.01) |
We compare P-DRUM against ensemble, MC-dropout [gal2016dropout], k-nearest neighbors of descriptors proposed by janet2019quantitative, and GMM of descriptors proposed by zhu2023fast. We used HME21 dataset which contains 37 elements [takamoto2022towards], and three datasets from rMD17 datasets [rmd17]: malondialdehyde, salicylic acid, uracil. Table 1 shows statistics of all datasets used in this paper. Additionally, we assess out-of-distribution (OOD) robustness with a benchmark on the nickel aluminide () dataset, generated using Matlantis [matlantis]. We obtained the MACE model from the official MACE repository 111https://github.com/ACEsuit/mace and used its implementation of the descriptor. Since MACE does not support MC-dropout by default, we implemented dropout in the fully connected layers after the activation function of both the interaction block and the readout block. Additional training details for our experiments are provided in Appendix B. In terms of computational efficiency, ensemble and MC-dropout require five forward passes of MACE, whereas kNN, GMM, and P-DRUM need only a single pass, with only negligible additional overhead compared to MACE forward pass cost.
4.1 Uncertainty-error correlation evaluation
Table 2 reports the 5-trial average Spearman correlation for each method across datasets. The results indicate that P-DRUM-diff performs relatively well in predicting energy uncertainties, whereas P-DRUM-norm is more effective for force uncertainties. Notably, the performance gap between GMM and kNN is larger in HME21. For readers who are interested, we provide an additional analysis using principal component analysis of HME21 in Appendix D.
4.2 OOD detection of
In this section, we evaluate the performance in the task of out-of-distribution (OOD) detection. We used the dataset, which was generated using Preferred Potential (PFP) [takamoto2022towards] within the Matlantis platform [matlantis], where the initial structure is collected from Materials Project (mp-2593) 222https://next-gen.materialsproject.org/materials/mp-2593 [jain10materials]. To assess OOD detection, we prepared several distinct OOD datasets for : (1) High-temperature OOD: structures generated via molecular dynamics simulations at temperatures higher than those used in the training dataset. The training data included temperatures of 500K, 1000K, and 1500K, while the OOD data was derived from simulations at 2000K and 3000K. (2) Hexagonal: with different phase from original dataset (mp-1183232).(3) Cubic: with different phase from the original dataset (mp-672232).(4) Swap: randomly swap positions of Ni and Al in the structures for 2, 4, and 8 pairs. We used PFP predictions as ground truths and force uncertainty for evaluation.
Table 3 summarizes the performance of various OOD detection methods. Ensemble, kNN, GMM, and P-DRUM-diff consistently achieve strong OOD detection performance across the benchmark. In comparison, P-DRUM-norm performs less effectively, while MC-dropout yields the lowest performance in terms of both Spearman correlation and AUC. Although P-DRUM-norm achieves the best performance in Table 2, it proves less effective in the OOD setting in our experiments.
| Method | High temp. | Hexagonal | Cubic | Swap | All | |||||
| Corr. | AUC | Corr. | AUC | Corr. | AUC | Corr. | AUC | Corr. | AUC | |
| Ensemble | 0.98 | 1.00 | 0.88 | 0.94 | 0.95 | 1.00 | 0.76 | 1.00 | 0.90 | 0.99 |
| MC-dropout [gal2016dropout] | 0.92 | 1.00 | 0.54 | 0.63 | 0.81 | 0.84 | 0.62 | 0.82 | 0.72 | 0.82 |
| GMM [zhu2023fast] | 0.98 | 1.00 | 0.74 | 1.00 | 0.73 | 1.00 | 0.77 | 1.00 | 0.81 | 1.00 |
| kNN [janet2019quantitative] | 0.98 | 1.00 | 0.80 | 0.99 | 0.75 | 1.00 | 0.76 | 1.00 | 0.82 | 1.00 |
| P-DRUM-norm | 0.99 | 1.00 | 0.82 | 0.82 | 0.78 | 0.82 | 0.70 | 0.99 | 0.82 | 0.91 |
| P-DRUM-diff | 0.97 | 1.00 | 0.89 | 0.97 | 0.80 | 1.00 | 0.81 | 1.00 | 0.87 | 0.99 |
5 Discussions
When P-DRUM outperform kNN, GMM? While descriptor-based methods perform competitively, P-DRUM clearly outperforms kNN and GMM in error–uncertainty correlation on the HME21 dataset. We hypothesize that when a dataset contains many elements, descriptor information alone may be difficult to capture the error correlation, and incorporating the prediction error signal can enhance the uncertainty estimation by aligning it more closely with the true prediction error.
Which approach is better: error norm learning or deviation learning? In Table 2, deviation learning outperforms error-norm learning for energy Spearman correlation, whereas the opposite trend is observed for force correlation. This difference likely arises from the nature of the targets: energy deviation is a scalar, where retaining the error sign aids learning, while force deviation is a three-dimensional vector, making direct estimation more challenging. Using the force error norm reduces this complexity, resulting in an improvement in the Spearman correlation. However, as shown in Table 3, using force uncertainty, error-norm learning underperforms deviation learning in OOD detection, indicating the need for further investigation into the advantages and limitations of these approaches. We hypothesize that compressing errors (-dimensional for forces) into a norm (-dimensional) might be detrimental for OOD detection.
6 Conclusion and future work
We investigated the effectiveness of post-hoc descriptor-based residual uncertainty modeling (P-DRUM) for machine learning interatomic potentials (MLIPs). P-DRUM achieved a higher Spearman correlation with prediction errors compared to other methods. However, P-DRUM-norm showed inferior out-of-distribution (OOD) detection performance compared to Gaussian mixture model (GMM) and k-nearest neighbor (kNN) descriptor-based approaches, revealing its potential limitations. Future work will explore extending P-DRUM to active learning pipelines and broader MLIP applications to further evaluate its versatility, as well as developing improved training strategies to enhance its performance. In addition, assessing the reliability of reusing the training set to construct the residual training set is an important direction. Using separate dataset splits for training the MLIP and the uncertainty estimation model can reduce bias from the original data but also limits the available training samples. Exploring this trade-off is useful to developing a data-efficient P-DRUM strategy.
Acknowledgments and Disclosure of Funding
We would like to thank Professor Ju Li for helpful comments and suggestions.
| Error type | Method | Uracil | Salicylic | Malondialdehyde | HME21 | |
| MC-dropout [gal2016dropout] | -0.48 (0.01) | 0.49 (0.02) | 0.50 (0.01) | 0.49 (0.04) | 0.55 (0.02) | |
| GMM [zhu2023fast] | 0.47 (0.06) | 0.46 (0.07) | 0.55 (0.04) | 0.71 (0.04) | 0.54 (0.01) | |
| kNN [janet2019quantitative] | 0.48 (0.05) | 0.47 (0.05) | 0.54 (0.03) | 0.70 (0.04) | 0.51 (0.01) | |
| P-DRUM-norm | 0.51 (0.03) | 0.48 (0.01) | 0.46 (0.02) | 0.72 (0.03) | 0.63 (0.02) | |
| P-DRUM-diff | 0.56 (0.07) | 0.57 (0.06) | 0.59 (0.06) | 0.90 (0.04) | 0.61 (0.01) | |
| MC-dropout [gal2016dropout] | 0.62 (0.03) | 0.63 (0.02) | 0.63 (0.04) | 0.95 (0.01) | 0.83 (0.01) | |
| GMM [zhu2023fast] | 0.79 (0.00) | 0.84 (0.02) | 0.82 (0.02) | 0.98 (0.01) | 0.85 (0.03) | |
| kNN [janet2019quantitative] | 0.77 (0.01) | 0.81 (0.02) | 0.81 (0.02) | 0.98 (0.01) | 0.84 (0.00) | |
| P-DRUM-norm | 0.85 (0.01) | 0.86 (0.03) | 0.83 (0.01) | 0.98 (0.00) | 0.98 (0.00) | |
| P-DRUM-diff | 0.76 (0.01) | 0.78 (0.02) | 0.77 (0.02) | 0.96 (0.06) | 0.92 (0.01) |
Appendix A Broader Impact
This work focuses on developing uncertainty estimation techniques called post-hoc descriptor-based residual uncertainty modeling (P-DRUM) for machine learning interatomic potentials. P-DRUM is designed to improve the reliability and robustness of simulations in chemistry, materials science, and related fields. Importantly, our research does not involve human subjects, personal data, or any form of unethical experimentation. The techniques we propose are purely computational and are evaluated on standard benchmark datasets and simulated molecular systems.
Appendix B Training details and hyperparameters